OpenStudy (tainted):

A compound that is composed of molybdenum (Mo) and oxygen (O) was produced in a lab by heating molybdenum over a Bunsen burner. The following data was collected: Mass of crucible: 38.26 g Mass of crucible and molybdenum: 39.52 g Mass of crucible and molybdenum oxide: 39.84 g Solve for the empirical formula of the compound, showing (or explaining in complete sentences) your calculations

OpenStudy (anonymous):

To get the mass of molybdenum in the crucible, you subtract the mass of the crucible from the mass of the crucible and molybdenum: 39.52g - 38.26g = 1.26g To get the mass of molybdenum oxide in the crucible, you subtract the mass of the crucible from the mass of the crucible and molybdenum oxide: 39.84g - 38.26g = 1.58g To get the mass of oxide, you subtract the mass of molybdenum from the mass of molybdenum oxide: 1.58g - 1.26g = 0.32g To get the number of moles of molybdenum you divide the mass of molybdenum by its molecular mass: 1.26g/96 = 0.01 moles To get the number of moles of oxide, divide the mass of oxide by its molecular mass: 0.32g/16 = 0.02 moles There are twice as many moles of oxygen so the empirical formula is: MoO2

OpenStudy (tainted):

Thanks!

OpenStudy (kira_yamato):

I used a slightly more precise value for the Ar and I got a different answer...

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS so using this answer will be an academic integrity violation. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065